# On the interpretation of quantum theory as games between observers and nature played in Minkowski spacetime

### v2

@@ 3.6.1 Causal contextuality scenario associated with an alternating spacetime game
Any alternating spacetime game can be associated with a causal contextuality scenario as follows. As explained in the previous tabular correspondence, the general idea of the overall mapping is that
@@
We must prove that the enabling relations have consistent left-hand sides for it to be a valid causal contextuality scenario.
@@
Note that $\mathcal{C}_{t}$ cannot be empty because there are no unused settings, and it also does not contain the empty set.
@@ 3.6.2 Alternating spacetime game associated with a causal contextuality scenario
    We also note that there cannot be any duplicates, i.e., there cannot be distinct nodes in $x$ with the same parent $t$ and indexed by the same $c\in\chi(t)$, because of BA2.
@@
We define causal contextuality scenarios that have a good cover999also more recently called causally secured covers by Abramsky et al. as the range of the mapping of Definition 7. Causal contextuality scenarios with a cover that is not a good cover can still be mapped to an alternating spacetime game, however, this mapping does not roundtrip and the correspondence has no semantic or physical meaning in this case.
@@
  • Map back to a causal contextuality scenario, which will have a good cover by definition of a good cover.
@@
3.7 The strategy sheaf of alternating games
Let us first consider the particular case of spacetime games with perfect information and the corresponding causal contextuality scenarios
If we take over the definition of a complete strategy of nature as a pure strategy appearing in the (reduced) strategic form (Fourny, 2020) of the alternating spacetime game—which is, indirectly, the (reduced) strategic form (Kuhn, 1950) of the associated game in extensive form with imperfect information—instead of that given in (Abramsky et al., 2024) in terms of game semantics, we note that these strategies are simply partial functions on a discrete space ($\mathcal{I}$) to actions. Thus, we generally consider the restrictions of these partial functions to smaller domains, from which it follows that the suchly defined presheaf of strategies has the sheaf property. This was left as an open question by Abramsky et al. (2024) to determine “good covers” for which this is the case, and we have thus shown that the natural cover of the associated game in this paper is such a good cover.
@@ 3.7 The strategy sheaf of alternating games
The argument why the sheaf property obtains is identical to the one given by Abramsky and Brandenburger (2011) for the event sheaf because it also involves the gluing of partial functions on discrete space that agree on their overlaps.
This result can naturally be extended to spacetime games with imperfect information: the conversion to the equivalent extensive form is made as for perfect information, and then, for each non-singleton information set in the spacetime game, the information sets corresponding, in the extensive form, to the individual nodes of the information set are merged.
Then, the pure strategies of nature for the game in extensive form are taken as a basis for the strategy sheaf.
@@
Finally, the simplicity of the construction of the strategy sheaf of Nature demonstrates the expressive power of the spacetime games framework, itself building on 70 years of game theory research.
3.8 Mixed strategies, Bell inequalities, and the strategy distribution presheaf
@@
Nash equilibria (Nash, 1951) and other concepts apply naturally to spacetime games (Fourny, 2020). In particular, a mixed strategy in the Nash sense directly corresponds to a deterministic hidden variable model where $\lambda$ is a pure strategy. Free choice, as commonly assumed in the quantum foundations’ literature, directly translates to nature and the observers independently picking their mixed strategies – as is done when considering Nash equilibria in classical game theory.
@@
If we observe that X and Y have disjoint supports (because Alice either picks basis x or y) and we merge them into a single one (call it X) – and we do the same for Bob (W and Z are merged into a single variable Z) – then the formula for computing the probability for the complete history resulting from a combination of a mixed strategy of nature (a distribution over the values of $\Lambda$) and the mixed strategies of observers (distributions over the values of $A$ and $B$) happens to exactly correspond to the formula used in the Bell inequality literature (Colbeck, 2017) for local-realist theories with measurement independence:
@@
Figure 10 shows how this can also be interpreted as a causal network, with all variables depending deterministically on the mixed strategies, leading to a distribution of the complete histories for each combination of mixed strategies of Alice, Bob, and Alfred. Since in our non-Nashian paradigm (Fourny, 2019a)(Baczyk and Fourny, 2023) we posit a symmetry between nature and observers (or equivalently, between settings and outcomes), we add explicit lambdas also for the observers and not only for nature.
@@
In the special case of flat games (no adaptive measurements), it was shown by Abramsky and Brandenburger (2011) that the existence of a deterministic hidden variable model realizing an empirical model is implied by the existence of a global section for it. We observe that a mixed strategy of nature is the same mathematical object as a global section of the event distribution presheaf and we remember that spacetime games have a natural cover given by the projection of the support of their complete histories to nature. Then, one must conclude that the existence of a mixed strategy of nature leading to a distribution on the complete histories that is consistent with an empirical model (on the natural cover) implies the existence of a deterministic hidden variable model for this empirical model. This is consistent with our qualitative argument given by Baczyk and Fourny (2023) that Nashian game theory is not compatible with quantum physics.
4 Discussion
@@ 4.1 Perfect information
Fourny (2020) observed that imperfect information in the extensive form of the game corresponding to a spacetime game with perfect information can be interpreted in terms of the preservation of locality: information sets are grouped in the way they are because of the finite speed of light, and spacelike separation.
If we do not require perfect information for spacetime games, this introduces another physical justification for imperfect information in the resulting extensive form, namely, noncontextuality, because non-singleton information sets in the spacetime game (which lead to the fusion of the corresponding information sets in the extensive form) arise from the different contexts in which a measurement can be carried out, with the result independent of the context. This is also consistent with the fact that locality is a special case of non-contextuality.
Having said this, we would like to conclude by sharing our scientific belief that only spacetime games with perfect information correspond to a physical reality, and that spacetime games with imperfect information are, above all, a useful mathematical tool for reductio ad absurdum proofs that nature is contextual. This is aligned with Dzhafarov et al. (2016) and on keeping locality as a fundamental principle of nature, provided we weaken measurement independence and assume non-Nashian free choice instead.
@@ 5 Related work
Hardy models (Hardy, 1993), PR boxes (Popescu and Rohrlich, 1994), and GHZ models (Greenberger et al., 1989) are representable with spacetime games.
@@
The framework also maps to contextuality by default (Dzhafarov et al., 2016), which corresponds to having perfect information. Forcing random variables corresponding to measurement outcomes to have the same value formally corresponds to grouping the corresponding nodes into the same information set.
@@
Sheaves of sections as constructed by (Gogioso and Pinzani, 2023) correspond to a special case of spacetime games with perfect information in which measurements are performed independently of previous measurement outcomes, without merging the information sets (see also Section 6.b. in (Abramsky et al., 2024)), and with unique causal bridges. The fact that the sheaf property obtains is consistent with our results, which generalize the case covered in (Gogioso and Pinzani, 2023) to all alternating spacetime games, and causal contextuality scenarios that have a natural cover.
@@
A definition of free choice is given by Renner and Colbeck (2011), as a decision (called spacetime variable) being independent of anything it could not have caused. A deterministic resolution algorithm for all spacetime games with perfect or imperfect information, and which circumvents impossibility theorems (such as the sheaf property of the strategy distribution presheaf not generally holding) by weakening free choice, is given in (Fourny, 2019a), with the special case of strategic games covered in (Fourny, 2017) and for games in extensive form with perfect information in (Fourny et al., 2014). A high-level discussion of the non-Nashian paradigm in the context of quantum physics interpreted as a game between the observer and nature is given in (Fourny, 2019b).
@@
The seeding philosophy of the non-Nashian approach is attributable to Dupuy (1992) (counterfactual decision theory) and the game theoretic seed and initial conjectures were made by Dupuy (2000).
@@
There is also subsequent work by Kadnikov and Wichardt (2022) (also see (Kadnikov, 2020)) that provides a strategic decision-making model generalizing the normal form and extensive form with graph structures related to spacetime games. This work defines the outcomes of the model based on sets of leaves of the graph structure and nicely connects it to the set-theoretic form by Bonanno (1992) (see also (Bonanno, 1991)). These sets of leaves can be put in correspondence with the complete histories, and the outcomes of the game in extensive form with imperfect information, giving additional and complementary insights.
@@
Hance and Hossenfelder (2022) give an excellent argument of why impossibility theorems do not disprove the existence of local theories. Hance et al. (2022) also is of the opinion that free choice can be weakened. Wharton and Argaman (2020) refers to this possibility as (non-)future-input-dependent parameters, and it is also known as measurement independence (see for example (Storz et al., 2023) for the explicit mention of this assumption in Bell experiments) or setting independence (Müller and Placek, 2023).
@@ 6 Acknowledgements
I thank my direct co-authors Jean-Pierre Dupuy, Stéphane Reiche, Michael Baczyk, Felipe Sulser, as well as my students at ETH who contributed to the game framework in various projects over the years: Emilien Pilloud, Ramon Gomm, Luc Stoffer, Florian Pauschitz. I also thank Michal Friedman, Richard Shiffrin, Renato Renner, V Vilasini, Philipp Kammerlander, Kuntal Sengupta, Ognyan Oreshkov, Tobias Sutter, Marek Szopa, Stefan Weigert, Beatrix Hiesmayr, Ben Moseley, Alexis Flesch, Lukas Sigrist, Philippe Colo, Abishek Ramdas, Günther Dissertori, Antoine Bommier, Gustavo Alonso, Daniel Kadnikov, Joseph Halpern, Robert French, Alexei Grinbaum, Matthew Jones, and VIS, for enlightening and stimulating discussions and for their support.

### v3

@@ 4.6.1 Causal contextuality scenario associated with an alternating spacetime game
Any alternating spacetime game can be associated with a causal contextuality scenario as follows. As explained in the previous tabular correspondence, the general idea of the overall mapping is that:
@@
For it to be a valid causal contextuality scenario, we must prove that the enabling relations have consistent left-hand sides.
@@
Note that $\mathcal{C}_{t}$ cannot be empty because there are no unused settings, and it does not contain the empty set.
@@ 4.6.2 Alternating spacetime game associated with a causal contextuality scenario
    We also note that there cannot be any duplicates, i.e., there cannot be distinct nodes in $x$ with the same parent $t$ and indexed by the same $c\in\chi(t)$ because of BA2.
@@
5 Concrete impact of the mapping on the simplicity of constructs and proofs
We now turn to a few examples showing that looking at a causal contextuality scenario via its associated spacetime game is directly useful and simplifies certain constructs, reasonings, and proofs.
5.1 Good covers
In the paper introducing causal contextuality scenarios [8], it is conjectured that certain covers dubbed “good covers” characterize certain causal contextuality scenarios for which desirable properties apply—such as the sheaf property. We first show that these good covers are straightforward to define from the mapping provided in this paper; in the next section, we will show that it is also straightforward to prove that the sheaf property applies in the case of pure strategies.
Concretely, causal contextuality scenarios that have a good cover can be defined as the range of the mapping of Definition 7. Causal contextuality scenarios with a cover that is not a good cover can still be mapped to an alternating spacetime game, however, this mapping does not roundtrip and the correspondence has no semantic or physical meaning in this case.
@@ 5.1 Good covers
It is also useful to know that one can change the cover of any causal contextuality scenario to be a good cover, keeping its other components intact.
@@
  • Map back to a causal contextuality scenario, which will have a good cover by construction.
@@
5.2 Global strategies
The paper introducing causal contextuality scenarios [8] defines strategies of nature and of the experimenter for a given causal contextuality scenario. Follow-up work [20] also mentions that these strategies can be expressed as global strategies that map measurements to outcomes, a construct also referred to—in our understanding— as the “flattening construction” in an IQSA keynote speech.
If we view a causal contextuality scenario as its corresponding spacetime game, this game has a corresponding strategic form as we explained in 2020 [10]: it is the (reduced) strategic form of the associated game in extensive form with imperfect information, which was constructed 75 years ago by Harold Kuhn [12][9]. A pure strategy in this strategic form provides the desired global strategy function.
5.3 The sheaf of strategies
Let us take any causal contextuality scenario with good cover.
@@ 5.3 The sheaf of strategies
In the flat case, it was shown by Abramsky and Brandenburger [21] that the presheaf of strategies has the sheaf property because it also involves the gluing of partial functions on discrete space that agree on their overlaps. Thanks to the spacetime game framework, it is straightforward to translate this argument to any causal contextuality scenario with good cover.
Indeed, if we take a pure strategy as defined by Kuhn instead of the form given by Abramsky et al [8] in terms of game semantics, we note that these strategies are partial functions on a discrete space ($\mathcal{I}$) to actions. Thus, we can generally consider the restrictions of these partial functions to smaller domains, from which it follows as in shown by Abramsky and Brandenburger [21] that the suchly defined presheaf of strategies has the sheaf property.
@@
This was left as an open question by [8] to determine “good covers” for which this is the case, and we have thus shown that Definition 9 successfully addresses this open question. The simplicity of the construction of the strategy sheaf of Nature demonstrates the expressive power of the spacetime game framework, itself building on 75 years of game theory literature.
5.4 Mixed strategies and empirical models
Nash equilibria [7] and other concepts apply naturally to spacetime games [10] through their associated extensive form. The strategic form constructed by Kuhn has pure strategies but also mixed strategies, which are probabilistic mixtures of pure strategies, i.e., a probability distribution over all pure strategies of a given player.
@@ 5.4 Mixed strategies and empirical models
A mixed strategy of nature thus corresponds to a deterministic hidden variable model where $\lambda$ is a pure strategy. Free choice, as commonly assumed in the quantum foundations’ literature, translates to nature and the observers independently picking their mixed strategies – as is done when considering Nash equilibria in classical game theory.
@@
Let us take the game in Figure 1 (Figure 2 for its extensive form with imperfect information) as an example. It corresponds to the 2-2-2 EPR experiment. If we observe that X and Y have disjoint supports (because Alice either picks basis x or y) and we merge them into a single one (call it X) – and we do the same for Bob (W and Z are merged into a single variable Z) – then the formula for computing the probability for the complete history resulting from a combination of a mixed strategy of nature (a distribution over the values of $\Lambda$) and the mixed strategies of observers (distributions over the values of $A$ and $B$) happens to exactly correspond to the formula used in the Bell inequality literature [22] for local-realist theories with measurement independence:
@@
Figure 10 shows how this can also be interpreted as a causal network, with all variables depending deterministically on the mixed strategies, leading to a distribution of the complete histories for each combination of mixed strategies of Alice, Bob, and Alfred. Since in our non-Nashian paradigm [5][6] we posit a symmetry between nature and observers (or equivalently, between settings and outcomes), we add explicit lambdas also for the observers and not only for nature.
@@
In the special case of flat games (no adaptive measurements), it was shown by Abramsky and Brandenburger [21] that the existence of a deterministic hidden variable model realizing an empirical model is implied by the existence of a global section for it. With spacetime games, this is intuitive to understand: we observe that a mixed strategy of nature is the same mathematical object as a global section of the event distribution presheaf and we remember that spacetime games have a natural cover given by the projection of the support of their complete histories to nature.
Then, the following result follows from Harold Kuhn’s strategic form:
Theorem 7.
Given a causal contextuality scenario with a good cover and an empirical model, the existence of a global section that is consistent with an empirical model implies the existence of a deterministic hidden variable model for this empirical model.
Proof.
If we consider the associated spacetime game, the existence of a global section consistent with the empirical model is equivalent to the existence of a mixed strategy of nature, i.e., a distribution on nature’s pure strategies, that is consistent with the empirical model. This mixed strategy provides a deterministic hidden variable model for this empirical model. ∎
We have just generalized the result of Abramsky and Brandenburger [21] from the flat case to any causal contextuality scenario with a good cover. The reason for the short proof is not that it is trivial; rather, it builds on previous work including Kuhn’s construct. This is consistent with our qualitative argument given by Baczyk and Fourny [6] that Nashian game theory is not compatible with quantum physics. The formulation in terms of presheaves and the sheaf property [21] strengthens this argument.
@@ 5.5 Perfect information
We observed [10] that imperfect information in the extensive form of the game corresponding to a spacetime game with perfect information can be interpreted in terms of the preservation of locality: information sets are grouped in the way they are because of the finite speed of light, and spacelike separation.
If we do not require perfect information for spacetime games, this introduces another physical justification for imperfect information in the resulting extensive form, namely, noncontextuality because non-singleton information sets in the spacetime game (which lead to the fusion of the corresponding information sets in the extensive form) arise from the different contexts in which a measurement can be carried out, with the result independent of the context. This is also consistent with the fact that locality is a special case of non-contextuality.
5.6 Temporal contextuality scenarios
Temporal contextuality scenarios were introduced by Searle et al [20] as a special case of causal contextuality scenario in which the measurements are timelike-separated. If we consider the associated spacetime games, temporal contextuality scenarios correspond to games in extensive form with perfect information. Furthermore, the lookbacks—corresponding to limited memory on past experiments— can be modeled in the spacetime game framework with imperfect information, i.e., by connecting identical measurements carried out in distinct pasts together. All the above simplified constructs apply: causal secured covers, global strategies, the presheaf of pure strategies, and the link between the existence of a global section of an empirical model and the existence of a deterministic hidden variable model.
@@ 6 Related work
Hardy models [14], PR boxes [15], and GHZ models [16] are representable with spacetime games.
@@
The framework also maps to contextuality by default [23], which corresponds to having perfect information. Forcing random variables corresponding to measurement outcomes to have the same value formally corresponds to grouping the corresponding nodes into the same information set.
@@
Sheaves of sections as constructed by Gogioso and Pinzani [13] correspond to a special case of spacetime games with perfect information in which measurements are performed independently of previous measurement outcomes, without merging the information sets (see also Section 6.b. in [8]), and with unique causal bridges. The fact that the sheaf property obtains is consistent with our results, which generalize the case covered by Gogioso and Pinzani [13] to all alternating spacetime games, and causal contextuality scenarios that have a natural cover.
@@
A definition of free choice is given by Renner and Colbeck [24], as a decision (called spacetime variable) being independent of anything it could not have caused. A deterministic resolution algorithm for all spacetime games with perfect or imperfect information, and which circumvents impossibility theorems (such as the sheaf property of the strategy distribution presheaf not generally holding) by weakening free choice, is given in [5], with the special case of games in normal form [25] and of games in extensive form with perfect information [17]. A high-level discussion of the non-Nashian paradigm in the context of quantum physics interpreted as a game between the observer and nature is also available [26].
@@
The seeding philosophy of the non-Nashian approach as well as the notion of causal bridge is attributable to Dupuy [18] (counterfactual decision theory) and the game theoretic seed and initial conjectures were made by [19].
@@
There is also subsequent work by Kadnikov and Wichardt [27][28] that provides a strategic decision-making model generalizing the normal form and extensive form with graph structures related to spacetime games. This work defines the outcomes of the model based on sets of leaves of the graph structure and nicely connects it to the set-theoretic form by Bonanno [29][30]. These sets of leaves can be put in correspondence with the complete histories, and the outcomes of the game in extensive form with imperfect information, giving additional and complementary insights.
@@
[31] give an excellent argument of why impossibility theorems do not disprove the existence of local theories. Hance and Hossenfelder [32] also are of the opinion that free choice can be weakened. Wharton and Argaman [33] refer to this possibility as (non-)future-input-dependent parameters, and it is also known as measurement independence (see for example the work of Storz et al [34] for the explicit mention of this assumption in Bell experiments) or setting independence [35].
7 Conclusion
We would like to conclude by sharing our scientific belief that only spacetime games with perfect information correspond to a physical reality, and that spacetime games with imperfect information are, above all, a useful mathematical tool for reductio ad absurdum proofs that nature is contextual. This is aligned with work by Dzhafarov [23] and on keeping locality as a fundamental principle of nature, provided we weaken measurement independence and assume non-Nashian free choice instead [6].
@@ 8 Acknowledgements
I thank my students at ETH who contributed to the game framework in various projects over the years: Felipe Sulser, Emilien Pilloud, Ramon Gomm, Luc Stoffer, Martin Kucera, Michael Baczyk, Florian Pauschitz.